OpenStudy (anonymous):

The sum of two numbers is 62 and their difference is 16. Which system of equations represents the word problem? 1. x + y = 62 and x - y = 16 2. x + y = 62 and 62-x = 16 3. x + 62 = y and x - 16 = y

OpenStudy (ajprincess):

Suppose the sum of two numbers is 6 and the difference is 2 then the system of equations that represent them are x+y=6-sum x-y=2-differnce Can u do nw?

OpenStudy (anonymous):

yes thank you.

OpenStudy (ajprincess):

yw.:)